OpenStudy (anonymous):

Find the slope and y-intercept of the line. y = –x + 7 A. slope: –7, y-intercept: 1 B. slope: –1, y-intercept: 7 C. slope: 1, y-intercept: –7 D. slope: 7, y-intercept: –1 @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

b

OpenStudy (cggurumanjunath):

y intercept is the constant !

OpenStudy (anonymous):

David is selling floral arrangements. Each arrangement uses 1 vase and 12 roses. Each vase costs David $2.00. Let C be the total cost of the arrangement and r be the cost of 1 rose. Which equation should David use to find the total cost of each arrangement? A. 12 = C + 2r B. C = 2r + 12 C. 12C = r + 12 D. C = 12r + 2

OpenStudy (cggurumanjunath):

y=-x+7

OpenStudy (cggurumanjunath):

7 is constant . so y intercept is +7

OpenStudy (luigi0210):

One question at a time.

OpenStudy (cggurumanjunath):

dy/dx=-1

OpenStudy (anonymous):

No i can post 2 questions per post

OpenStudy (cggurumanjunath):

so slope is -1 .

OpenStudy (anonymous):

What about the second one?

OpenStudy (cggurumanjunath):

d